# Dense Video Object Captioning from Disjoint Supervision

### v2

@@ 4.5 Analysis of disjoint training
Finetuning evaluation. We now finetune each of the pretrained models from Tab. References and show results in Tab. References. We also include a baseline (#0) which initializes from only a CLIP-pretrained checkpoint [48], observing that this model performs poorly. If we pretrain on only a single dataset, we find that Visual Genome (VG) is the best, probably because it allows us to pretrain both the captioning- and object proposal modules (Fig. 3). Once again, we observe that different pretraining datasets are complementary, as adding either SMiT or COCO (#2 to #4,  #2 to #5,  #1 to #6) improves results further. Adding more pretraining datasets improves results further (#7), and we achieve the best results with our model pretrained on all pretraining datasets (#8), which outperforms the best single-dataset pretrained model by 2.0 CHOTA on VidSTG, and 0.7 CHOTA on VLN. The improvement over only a CLIP-pretrained checkpoint is even larger, by 9.1 CHOTA and 11.6 CHOTA on the two respective datasets. Qualitative visualizations are in Fig. 4 and the supplement.
@@
4.6 State-of-the-art comparison on video grounding
As introduced in Sec. References, models trained for Dense VOC can be directly used for sentence grounding, by finding the proposals with the maximum likelihood of generating the query. We evaluate spatial grounding on the VLN Location-QA [60] and VidSTG [83] benchmarks respectively.
VLN Location-QA consists of questions starting with “Where is”, and requires the model to produce a bounding box at each frame in the video. The task is therefore effectively a sentence grounding problem, and indeed, the ReferFormer [66] baseline used by [60] performs sentence grounding after removing “Where is” from the question. We also remove this prefix before grounding following Sec. References for both our final model, and an additional GRiT baseline.
In this dataset, only one annotated frame (unknown at inference time) is evaluated, and this benchmark therefore effectively does not involve temporal localization. As the annotation of this dataset is based on mouse traces instead of bounding boxes, the evaluation metric considers bounding box coverage (recall) and precision (full details in [60]). As shown in Tab. 6, we improve substantially over ReferFormer and our GRiT [65] baseline.
Table 5: State-of-the-art comparison of spatial grounding on VLN Location-QA [60]. We report the official evaluation metric [60], which evaluates if both bounding box recall and precision are above 0.5. We compare to the VLN ReferFormer baseline [60, 66], GRiT [65], and our full model (#8 of Tab. References).
Backbone Recall&Prec.$>0.5$
ReferFormer [66, 60] ResNet50 48.3
GRiT [65] ViT-B 62.1
Ours ViT-B 65.1
Finetuned Zero-shot
STVGBert [56] 47.3 -
TubeDETR [72] 59.0 -
STCAT [31] 61.7 -
Ours 61.9 54.1
Table 5: State-of-the-art comparison of spatial grounding on VLN Location-QA [60]. We report the official evaluation metric [60], which evaluates if both bounding box recall and precision are above 0.5. We compare to the VLN ReferFormer baseline [60, 66], GRiT [65], and our full model (#8 of Tab. References).
Table 6: State-of-the-art comparison of spatial grounding on the VidSTG test set. All models assume ground truth temporal localization is provided.
VidSTG requires producing a sequence of bounding boxes for a given sentence query. The evaluation metric is the average of the Intersection over Union (IoU) at each frame, between the predicted and ground truth bounding boxes for the target object. We compare to other prior works on this dataset in Tab. References, assuming that the input video is already trimmed temporally to the objects of interest. Our model achieves the best IoU, outperforming models designed specifically for grounding, thereby showing that our generative framework can be used effectively in the discriminative grounding task. In contrast to prior work, we also evaluate zero-shot without training on VidSTG, and still perform competitively. This result emphasizes the efficacy of our disjoint pretraining. We provide more results on VidSTG in the supplement.
5 Conclusion and Future Work
We proposed the new task of dense video object captioning. Although this task requires expensive annotations across space, time and language, we show that we can train a model on existing larger-scale datasets for disjoint subtasks. Moreover, our proposed end-to-end architecture is important for producing more accurate and coherent captions. Currently, our model produces a single caption for each trajectory, and in future work, we aim to caption potentially multiple action segments within one trajectory. We repurposed grounding datasets for our task, we aim to obtain a dataset with richer spatio-temporal captions in future.
Acknowledgments. We thank Shyamal Buch and Shen Yan for helpful feedback.
Appendix 0.A Additional Experimental and Implementation Details
0.A.1 $\text{AP}_{M}$implementation details
mAP-METEOR is the official evaluation metric used in Visual Genome [34] dataset for dense image object captioning. This metric evaluates predictions in each frame separately, without evaluating the tracking output.
mAP-METEOR is based on the Average Precision used in object detection [41, 24], but includes a caption similarity criteria for determining true positives: i.e. a prediction is a true positive if the Intersection over Union (IoU) with the ground truth bounding box is above a threshold, and if the METEOR score [5] is above another threshold. We follow the same implementation and thresholds as the Visual Genome dataset***https://github.com/jcjohnson/densecap/blob/master/eval/eval_utils.lua. i.e. IoU thresholds of (0.3, 0.4, 0.5, 0.6, 0.7) and METEOR thresholds of (0.0, 0.05, 0.1, 0.15, 0.2).
In our case, some objects in the datasets [83, 60] only have bounding box annotations and no caption annotations (Sec. References). For these objects without caption annotations, we allow any caption prediction (and therefore ignore it) by setting its METEOR score to the maximum of 1. For brevity, we abbreviated this metric as $\text{AP}_{M}$in the main paper.
0.A.2 CHOTA Details and Evaluation code
@@
We attach our CHOTA evaluation code as chota.py. This evaluation code is based on the official HOTA implementation***https://github.com/JonathonLuiten/TrackEval. The original code is under an MIT license.
0.A.3 Full training details
@@
As mentioned in Sec. References, our model is based on GRiT [65]. The original GRiT code***https://github.com/JialianW/GRiT is released under an MIT license. Following GRiT, we use a ViTDet-Base [22, 39] backbone, a CenterNet [88] region proposal network and RoI Head, and a randomly-initialized text decoder following that of GIT [61]. The text decoder consists of 6 self-attention layers with casual feature masks [61]. All model architecture parameters follow the defaults from GRiT [65].
@@
The original GRiT uses an MAE pretrained checkpoint, while in our case we found a CLIP pretrained checkpoint [48] performs better on our task. To fit more frames into memory for both training and evaluation, we use a $384\!\times\!384$ input size instead of the original $1024\!\times\!1024$. This choice moderately decreases dense image captioning performance on Visual Genome (from 17.3 $\text{AP}_{M}$to 15.7 $\text{AP}_{M}$).
@@
During disjoint multi-dataset pretraining, we sample batches from different datasets in an even ratio ($1:1:1:1$). For image datasets, a batch is composed of different images; for video datasets, we put the time dimension in batches and always guarantee images in the same mini-batch are from the same video. We use a local batch size of either 1 video (consisting of 8 sampled frames), or 8 images. As we use 32 TPUs, this means that our global batch size is either 32 videos or 256 images. We use the AdamW optimizer with a learning rate of $2\times 10^{-4}$, weight decay of $0.05$, and a layerwise learning rate decay of $0.7$ [39, 65]. We train for $22.5\times 10^{3}$ iterations per dataset, decreasing the learning rate by a factor of 10 after $90\%$ and $97.5\%$ of the training schedule [65]. For pretraining on all the 4 datasets in Sec. References, this corresponds to a total of $90\times 10^{3}$ iterations, which took approximately 20 hours on 32, 16GB TPUs cores.
@@
For VidSTG [83] finetuning, we sample 16 frames in training, and run on all 200 frames in testing. For VLN [60] finetuning, we use the 3 annotated frames in both training and evaluation. For finetuning experiments on both datasets, we use a video batch size $16$ and train for $11.25\times 10^{3}$ iterations, with a learning rate of $10^{-5}$, weight decay of $0.05$, and layerwise-learning decay of $0.7$ [39]. The finetuning took approximately 6 hours on 16, 16GB TPUs for VidSTG, and about 2 hours on 16, 16GB TPUs for VLN. Inference on VidSTG requires 32GB of TPU memory to fit 200 frames.
@@
The detection loss [88] involves a center heatmap loss, a bounding box regression loss, and a classification and bounding box refinement loss in the RoI head:
@@
$\scriptsize L_{heatmap}(Y,\bar{Y})=\frac{1}{n}\sum_{ij}\begin{cases}(1-{Y}_{ij})^{\alpha}\log({Y}_{ij})&\!\text{if}\ \bar{Y}_{ij}=1\\ (1-\bar{Y}_{ij})^{\beta}({Y}_{ij})^{\alpha}\log(1-{Y}_{ij})&\!\text{otherwise,}\end{cases}$ (5)
where $n$ is the number of objects in the image, $\alpha=2$ and $\beta=4$ are the focal loss weights [40].
$L_{reg}$ is a gIoU loss [58]:
$\footnotesize L_{reg}(B,\bar{B})=\frac{1}{n}\sum_{i}(\text{IoU}(B_{i},\bar{B}_{i})-\frac{|C_{i}\backslash(B_{i}\cup\bar{B}_{i})|}{|C_{i}|}),$ (6)
@@
$\footnotesize L_{assoc}=-\frac{1}{M}\sum_{ij}(\hat{A}_{ij}\log{A_{ij}}+(1-\hat{A}_{ij})\log{(1-A_{ij})})).$ (9)
The captioning loss is a softmax on each predicted word over the entire vocabulary, with a label smoothing co-efficient of $0.1$ following GIT [61].
@@
Appendix 0.B Additional Experimental Analysis
0.B.1 Ablation of hard tracking sampling.
We analyze the effect of the number of sampled frames, $m$, in hard-aggregation (Sec. References) in Tab. 7. With hard-aggregation, the captioning accuracy benefits from a larger number of frames $m$, thanks to longer input-sequence length. However, this also costs more TPU memory in both training and testing. We use $m=6$ in our ablation experiments (Tab. References) as it achieves the best accuracy. It also follows the default number of frames used in the GIT [61] video captioning model.
@@ Appendix 0.B Additional Experimental Analysis
Table 7: Hyper-parameter sweep for number of sampled frames, $m$, for hard-tracking. We show results on VidSTG [83] validation set. The models are based on #2 of Tab. References on VidSTG. Results with hard-feature aggregation get improved with more sampled frames and get saturated with 6 frames.
@@
Table 8: Results using UVO [63] as the tracking dataset. We show both zero-shot results (top) and finetuning results (bottom) on VidSTG datasets. For reference, we also include our results of using Aug-COCO (#8 of Tab. References and Tab. References). Aug-COCO performs better in both settings, motivating our choice in the main paper.
0.B.2 Using the UVO dataset for disjoint pretraining
For the disjoint pretraining of our model (Sec. References), we used Augmented COCO as our tracking dataset. Another alternative would have been to use UVO [63], which contains real-world videos, but is relatively small at only 5000 videos.
Table 8 compares Aug-COCO and UVO under the setting of Tab. References and Tab. References #8, both using a default multi-dataset sampling ratio $1:1:1:1$. We observe that disjoint pretraining with Aug-COCO consistently performs better than UVO in both zero-shot and finetuning scenarios, thus motivating our choice to use Aug-COCO for our experiments in the main paper.
0.B.3 Detailed captioning results
The Captioning Accuracy (CapA) component of our CHOTA metric is the average of the CIDEr, METEOR and SPICE metrics. For completeness, we report each of these captioning metrics individually in Tabs. 9 and 10, for zero-shot and full-finetuning evaluation, respectively.

### v3

@@ 4.5 Analysis of disjoint training
Finetuning evaluation. We now finetune each of the pretrained models from Tab. 3 left and show results in the right. We also include a baseline (#0) which initializes from only a CLIP-pretrained checkpoint, observing that this model performs poorly. Once again, we observe that different pretraining datasets are complementary, as adding either SMiT or COCO (#2 to #4,  #2 to #5,  #1 to #6) improves results further. Adding more pretraining datasets improves results further (#7), and we achieve the best results with our model pretrained on all pretraining datasets (#8), which outperforms the best single-dataset pretrained model by 2.0 CHOTA on VidSTG, and 0.7 CHOTA on VLN. The improvement over only a CLIP-pretrained checkpoint is even larger, by 9.1 CHOTA and 11.6 CHOTA on the two respective datasets. Qualitative visualizations are shown in the supplement.
4.6 Comparison to concurrent works BenSMOT and OW-VISCap.
We compare to the concurrent work Li et al. (2024b), which focuses on person category rather than all classes like our task. We finetune our model on the training set using the same hyper-parameters as our VidSTG experiments. Our full model achieved 90.19 HOTA and 0.254 CIDEr on this benchmark, significantly outperforming Li et al. (2024b). There are two major advantages of our model: our disjoint pretraining, and our use of a larger backbone (ViT-B vs. DLA-34 in BeyondMOT). We further break down the improvements by removing these two components in Tab. 5. The results show: 1. Our pretraining provides consistent gains on Li et al. (2024b) benchmark, improving especially captioning metrics. 2. With a small backbone and no pretraining, our model still outperforms Li et al. (2024b) on tracking and captioning metrics, showing the advantages of our end-to-end architecture.
HOTA DetA AssA CIDEr
Li et al. (2024b) 71.98 80.79 73.71 0.087
Ours 90.19 90.79 89.59 0.254
    - Pretrain 88.56 89.38 87.74 0.150
      - Backbone 86.55 84.33 89.19 0.129
Table 4: State-of-the-art comparison on BenSMOT (Li et al., 2024b). Our model outperforms  Li et al. (2024b) on comparable setting (no extra data, small backbone), and our full model improved $18.2$ HOTA and 0.167 CIDEr.
CHOTA DetA AssA CapA
OW-VisCap 53.0 60.1 54.0 43.9
Ours 56.9 65.8 70.4 39.7
Table 5: Compare with concurrent work OW-VisCap (Choudhuri et al., 2024) on VidSTG. The results are from  Choudhuri et al. (2024) Tab. 2 and our Tab. 3 #8 under the same setting. Our model are better at detection and tracking, with lower captioning accuracy due to a smaller langauge head (46M vs. 2.7B params.).
We compare to OW-VISCap which uses a Mask2Former architecture with video object queries. Tab. 5 shows an improved overall performance in CHOTA. Our largest improvement is in Association Accuracy, showing that our end-to-end tracking module (Sec. 3.2) outperforms the Mask2Former counterparts. OW-VisCap gets a higher captioning accuracy as they used a substantially larger, 2.7 billion parameter OPT language decoder (Zhang et al., 2022a), whilst we used a smaller, 46 million parameter language model as in GIT (Wang et al., 2022).
4.7 State-of-the-art comparison on video grounding
As introduced in Sec. 3.5, Dense VOC models can be directly used for sentence grounding, by finding the proposals with the maximum likelihood of generating the query. We evaluate spatial grounding on the VLN Location-QA (Voigtlaender et al., 2023) and VidSTG (Zhang et al., 2020) benchmarks.
VLN Location-QA consists of questions starting with “Where is”, and requires the model to produce a bounding box at each frame in the video. The task is therefore effectively a sentence grounding problem, and indeed, the ReferFormer (Wu et al., 2022b) baseline used by Voigtlaender et al. (2023) performs sentence grounding after removing “Where is” from the question. We also remove this prefix before grounding following Sec. 3.5 for both our final model, and an additional GRiT baseline.
In this dataset, only one annotated frame (unknown at inference time) is evaluated, and this benchmark therefore effectively does not involve temporal localization. As the annotation of this dataset is based on mouse traces instead of bounding boxes, the evaluation metric considers bounding box coverage (recall) and precision (full details in Voigtlaender et al. (2023)). As shown in Tab. 7, we improve substantially over ReferFormer and our GRiT (Wu et al., 2022a) baseline.
Backbone Rec. & Prec.$>0.5$
ReferFormer ResNet50 48.3
GRiT ViT-B 62.1
Ours ViT-B 65.1
Table 6: State-of-the-art comparison of spatial grounding on VLN Location-QA Voigtlaender et al. (2023). We report the official metric, which evaluates if bounding box recall and precision are both above 0.5. We compare to the ReferFormer baseline Voigtlaender et al. (2023), GRiT Wu et al. (2022a), and our model (#8 of Tab. 3 left).
Finetuned Zero-shot
STVGBert 47.3 -
TubeDETR 59.0 -
STCAT 61.7 -
Ours 61.9 54.1
Table 7: State-of-the-art comparison of spatial grounding on the VidSTG with STVGBert Su et al. (2021), TubeDETR Yang et al. (2022), and STCAT Jin et al. (2022). All models use ground truth temporal localization.
VidSTG requires producing a sequence of bounding boxes for a given sentence query. The evaluation metric is the average of the Intersection over Union (IoU) at each frame, between the predicted and ground truth bounding boxes for the target object. We compare to other prior works on this dataset in Tab. 7, assuming that the input video is already trimmed temporally to the objects of interest. Our model achieves the best IoU, outperforming models designed specifically for grounding, thereby showing that our generative framework can be used effectively in the discriminative grounding task. We also evaluate zero-shot without training on VidSTG, and still perform competitively. This emphasizes the efficacy of our disjoint pretraining. We provide more results in Appendix C.5.
5 Conclusion
We proposed the new task of dense video object captioning. Although this task requires expensive annotations across space, time and language, we show that we can train a model on existing larger-scale datasets for disjoint subtasks. We show our proposed end-to-end architecture is important for producing more accurate and coherent captions.
Appendices
We present further qualitative results (App. A), additional experimental details (App. B), additional experimental analysis (App. C), and broader impact and potential negative impact (App. E).
Appendix A Qualitative Results
@@
We show example qualitative visualizations Fig. 4 and discuss typical failure cases.
Appendix B Additional Experimental and Implementation Details
B.1 Code
Our code is available at https://github.com/google-research/scenic.
@@ B.1 Code
Our CHOTA evaluation code is in file “code/chota.py”. This evaluation code is based on the official HOTA implementation***https://github.com/JonathonLuiten/TrackEval. The original code is under an MIT license.
B.2 Full training details
@@
As mentioned in Sec. 4.3, our model is based on GRiT Wu et al. (2022a). The original GRiT code***https://github.com/JialianW/GRiT is released under an MIT license. Following GRiT, we use a ViTDet-Base Dosovitskiy et al. (2021); Li et al. (2022b) backbone, a CenterNet Zhou et al. (2019) region proposal network and RoI Head, and a randomly-initialized text decoder following that of GIT Wang et al. (2022). The text decoder consists of 6 self-attention layers with casual feature masks Wang et al. (2022). All model architecture parameters follow the defaults from GRiT Wu et al. (2022a).
@@
The original GRiT uses an MAE pretrained checkpoint, while in our case we found a CLIP pretrained checkpoint Radford et al. (2021) performs better on our task. To fit more frames into memory for both training and evaluation, we use a $384\!\times\!384$ input size instead of the original $1024\!\times\!1024$. This choice moderately decreases dense image captioning performance on Visual Genome (from 17.3 $\text{AP}_{M}$to 15.7 $\text{AP}_{M}$).
@@
During disjoint multi-dataset pretraining, we sample batches from different datasets in an even ratio ($1:1:1:1$). For image datasets, a batch is composed of different images; for video datasets, we put the time dimension in batches and always guarantee images in the same mini-batch are from the same video. We use a local batch size of either 1 video (consisting of 8 sampled frames), or 8 images. As we use 32 GPUs, this means that our global batch size is either 32 videos or 256 images. We use the AdamW optimizer with a learning rate of $2\times 10^{-4}$, weight decay of $0.05$, and a layerwise learning rate decay of $0.7$ Li et al. (2022b); Wu et al. (2022a). We train for $22.5\times 10^{3}$ iterations per dataset, decreasing the learning rate by a factor of 10 after $90\%$ and $97.5\%$ of the training schedule Wu et al. (2022a). For pretraining on all the 4 datasets in Sec. 3.4, this corresponds to a total of $90\times 10^{3}$ iterations, which took approximately 20 hours on 32, 16GB V100 GPUs.
@@
For VidSTG Zhang et al. (2020) finetuning, we sample 16 frames in training, and run on all 200 frames in testing. For VLN Voigtlaender et al. (2023) finetuning, we use the 3 annotated frames in both training and evaluation. For finetuning experiments on both datasets, we use a video batch size $16$ and train for $11.25\times 10^{3}$ iterations, with a learning rate of $10^{-5}$, weight decay of $0.05$, and layerwise-learning decay of $0.7$ Li et al. (2022b). The finetuning took approximately 6 hours on 16, 16GB GPUs for VidSTG, and about 2 hours on 16, 16GB GPUs for VLN. Inference on VidSTG requires 32GB of GPU memory to fit 200 frames.
@@
The detection loss Zhou et al. (2019) involves a center heatmap loss, a bounding box regression loss, and a classification and bounding box refinement loss in the RoI head:
@@
$L_{heatmap}(Y,\bar{Y})=\frac{1}{n}\sum_{ij}\begin{cases}(1-{Y}_{ij})^{\alpha}\log({Y}_{ij})&\!\text{if}\ \bar{Y}_{ij}=1\vspace{2mm}\\ (1-\bar{Y}_{ij})^{\beta}({Y}_{ij})^{\alpha}\log(1-{Y}_{ij})&\!\text{otherwise,}\end{cases}$ (5)
where $n$ is the number of objects in the image, $\alpha=2$ and $\beta=4$ are the focal loss weights Lin et al. (2017).
$L_{reg}$ is a gIoU loss Union (2019):
$L_{reg}(B,\bar{B})=\frac{1}{n}\sum_{i}(\text{IoU}(B_{i},\bar{B}_{i})-\frac{|C_{i}\backslash(B_{i}\cup\bar{B}_{i})|}{|C_{i}|}),$ (6)
@@
$L_{assoc}=-\frac{1}{M}\sum_{ij}(\hat{A}_{ij}\log{A_{ij}}+(1-\hat{A}_{ij})\log{(1-A_{ij})})).$ (9)
The captioning loss is a softmax on each predicted word over the entire vocabulary, with a label smoothing co-efficient of $0.1$ following GIT Wang et al. (2022).
@@
Appendix C Additional Experimental Analysis
C.1 $\text{AP}_{M}$evaluation.
# COCO VG SMiT Aug VidSTG (zero-shot) VLN (zero-shot) VidSTG (finetuned) VLN (finetuned)
COCO $\text{AP}_{M}$ $\text{AP}_{M}$ $\text{AP}_{M}$ $\text{AP}_{M}$
0 - - 54.1 35.1
1 - - 69.1 36.3
2 17.1 9.9 68.7 45.9
3 - - 54.8 38.0
4 18.2 12.7 68.9 47.2
5 37.4 19.7 70.8 46.1
6 36.7 18.1 69.4 41.3
7 38.2 19.0 71.2 48.2
8 39.5 20.1 71.5 48.2
Table 8: Zero-shot (left) and finetuning (right) evaluation of our disjoint trained models with varying datasets using image dense captioning metric $\text{AP}_{M}$. We show results on VidSTG Zhang et al. (2020) and VLN Voigtlaender et al. (2023). Each row is a model pretrained on the specified datasets for zero-shot evaluation and then finetuned on the downstream datasets following Tab. 3. The results are consistent with the CHOTA metric: our models trained on joint datasets perform the best.
mAP-METEOR is the official evaluation metric used in Visual Genome Krishna et al. (2017b) dataset for dense image object captioning. This metric evaluates predictions in each frame separately, without evaluating the tracking output.
mAP-METEOR is based on the Average Precision used in object detection Lin et al. (2014); Everingham et al. (2015), but includes a caption similarity criteria for determining true positives: i.e. a prediction is a true positive if the Intersection over Union (IoU) with the ground truth bounding box is above a threshold, and if the METEOR score Banerjee & Lavie (2005) is above another threshold. We follow the same implementation and thresholds as the Visual Genome dataset***https://github.com/jcjohnson/densecap/blob/master/eval/eval_utils.lua. i.e. IoU thresholds of (0.3, 0.4, 0.5, 0.6, 0.7) and METEOR thresholds of (0.0, 0.05, 0.1, 0.15, 0.2).
In our case, some objects in the datasets Zhang et al. (2020); Voigtlaender et al. (2023) only have bounding box annotations and no caption annotations (Sec. 4.1). For these objects without caption annotations, we allow any caption prediction (and therefore ignore it) by setting its METEOR score to the maximum of 1. For brevity, we abbreviated this metric as $\text{AP}_{M}$. We report $\text{AP}_{M}$following Tab. 3 in Tab. 8. The improvements are consistent with CHOTA.
C.2 Ablation of hard tracking sampling.
We analyze the effect of the number of sampled frames, $m$, in hard-aggregation (Sec. 3.3) in Tab. 9. With hard-aggregation, the captioning accuracy benefits from a larger number of frames $m$, thanks to longer input-sequence length. However, this also costs more GPU memory in both training and testing. We use $m=6$ in our ablation experiments (Tab. 2) as it achieves the best accuracy. It also follows the default number of frames used in the GIT Wang et al. (2022) video captioning model.
@@ Appendix C Additional Experimental Analysis
Table 9: Hyper-parameter sweep for number of sampled frames, $m$, for hard-tracking. We show results on VidSTG Zhang et al. (2020) validation. The models are based on #2 of Tab. 3 right on VidSTG. Results with hard-feature aggregation get improved with more frames and get saturated with 6 frames.
@@
Table 10: Results using UVO Wang et al. (2021b) as the tracking dataset. We show both zero-shot results (top) and finetuning results (bottom) on VidSTG datasets. For reference, we also include our results of using Aug-COCO (#8 of Tab. 3). Aug-COCO performs better in both settings, motivating our choice.
C.3 Using the UVO dataset for disjoint pretraining
For the disjoint pretraining of our model (Sec. 3.4), we used Augmented COCO as our tracking dataset. Another alternative would have been to use UVO Wang et al. (2021b), which contains real-world videos, but is relatively small at only 5000 videos.
Table 10 compares Aug-COCO and UVO under the setting of Tab. 3 #8, both using a default multi-dataset sampling ratio $1:1:1:1$. We observe that disjoint pretraining with Aug-COCO consistently performs better than UVO in both zero-shot and finetuning scenarios, thus motivating our choice to use Aug-COCO for our experiments in the main paper.
C.4 Detailed captioning results
The Captioning Accuracy (CapA) component of our CHOTA metric is the average of the CIDEr, METEOR and SPICE metrics. For completeness, we report each of these captioning metrics individually in Tabs. 11 and 12, for zero-shot and full-finetuning evaluation, respectively.